Minimize f(x,y) = x² + y² subject to 2x + 4y = 20. x=
Answer #1

20 Solution f(x,y)= x+y glx,y): 9x+4y= =>9x¢bly a90=0 Lla,y) = f(x,y) + X. g(x,y) x+y+ 2(2x + y -20) f (x + yrit. 2 (2x + 4yI solved the problem step by step.Please feel free to comment if you have doubt in any step.Please up vote if you understand the solution. Thank you.